OpenStudy (anonymous):

SImplify by removing factors of 1 z^2+16/z+4

OpenStudy (saifoo.khan):

\[\frac{(z-4)(z+4)}{z+4}\]

OpenStudy (anonymous):

find all numbers for which the rational expression is undefined y-3/9

OpenStudy (saifoo.khan):

incomplete question

OpenStudy (anonymous):

thats all it says, thats why i dont understand it!

OpenStudy (saifoo.khan):

\[\frac{y-3}{9}\]?

OpenStudy (anonymous):

yes and it says find all numbers which the rational expression is undefined

OpenStudy (saifoo.khan):

it can be 3.

OpenStudy (anonymous):

did you see the newist one i posted?

OpenStudy (saifoo.khan):

Yes..

OpenStudy (anonymous):

any ideas?

OpenStudy (saifoo.khan):

It's answer is, \[\large \frac{64u^2z^5}{5}\]

OpenStudy (saifoo.khan):

Which week are u in?

OpenStudy (anonymous):

im not sure its an online class but im on exercise 12.1 if that helps any..

OpenStudy (saifoo.khan):

It has same questions. ;)

OpenStudy (anonymous):

7-y/y-7simplify by removing factors of 1?

OpenStudy (anonymous):

and what do you mean? what has the same questions?

OpenStudy (saifoo.khan):

\[\frac{7-y}{y-7} \to \frac{7-y}{-1(7-y)} = \frac{1}{-1} = -1\]